OpenStudy (anonymous):

find all solutions of the trigonometric equation 2cos(x) + tan(x) = sec(x)

OpenStudy (evanhelps):

So first thing you want to do is simplify everything into sine and cosine: 2cos(x) + sin(x)/cos(x) = 1/cos(x) Then simplify down: 2cos^2(x) + sin(x) = 1 A trigonometric identity states that 2cos^2(x) = 1 - sin^2(x) so: 1 - sin^2(x) + sin(x) = 1 -sin^2(x) + sin(x) = 0 Then you can substitute all terms sin(x) as a letter and solve as a quadratic. -u^2 + u = 0 And solve the quadratic to get solutions (0, 1)

OpenStudy (anonymous):

Oh, thanks!

OpenStudy (anonymous):

Then do I plug the solutions back into the original equation?

OpenStudy (evanhelps):

Yeah to test it.

OpenStudy (evanhelps):

But I might be wrong if I messed up an identity or something

OpenStudy (evanhelps):

Hold on I'll check it

OpenStudy (anonymous):

Okay

OpenStudy (evanhelps):

Yeah I did something wrong. Sorry I don't really remember trig identities that well

OpenStudy (evanhelps):

Do you have a sheet with them on it?

OpenStudy (anonymous):

The identities?

OpenStudy (anonymous):

Cos^2(x) + sin^2(x) = 1

OpenStudy (anonymous):

1+tan^2(x)=sec^2(x) 1+cot^2(x)=csc^2(x)

OpenStudy (evanhelps):

Sorry man I can't figure this one out

OpenStudy (anonymous):

It's good thanks anyways